# Bundle Neural Networks for message diffusion on graphs

### v1

@@ Appendix C Proofs
Defining the ‘universal bundle’: We fix $k=\sum_{\mathsf{G}\in\mathcal{K}}\left|\mathsf{V}_{\mathsf{G}}\right|$, so we can interpret our embedding space as a lookup table where each index corresponds to a node $v\in\bigsqcup_{\mathsf{G}\in\mathcal{K}}V_{\mathsf{G}}$. In turn, we can think of the parameter matrix $\mathbf{W}\in\mathbb{R}^{\left(\sum_{G\in\mathcal{K}}\left|\mathsf{V}_{\mathsf{G}}\right|\right)\times\left(\sum_{G\in\mathcal{K}}\left|\mathsf{V}_{\mathsf{G}}\right|\right)}$ as a lookup table where each entry corresponds to a pair of nodes in our dataset $\mathcal{K}$. Still thinking of the indices of the $2ck$ dimensions as $2c$-dimensional vectors indexed by the $k$ nodes in our dataset, we define $\mathbf{O}_{u}\in O\left(2ck\right)$ as a block diagonal matrix with $k$ different $2c$-dimensional blocks. These are all set to the identity except for the block at the index corresponding to node $u$, which is defined as $\begin{pmatrix}\mathbf{\mathbf{0}_{c\times c}}&\mathbf{I}_{c\times c}\\ \mathbf{I}_{c\times c}&\mathbf{\mathbf{0}_{c\times c}}\end{pmatrix}$ which is a $2c\times 2c$ matrix that acts by permuting the first $c$ dimensions with the second $c$ dimensions.
@@
$\begin{split}&\frac{\partial\left(\operatorname{BuNN}\left(\mathbf{X}\right)\right)_{u}}{\partial\mathbf{x}_{v}}\\ &=\operatorname{pool}\frac{\partial\left(\widehat{\operatorname{BuNN}}\left(\operatorname{lift}\left(\mathbf{X}\right)\right)\right)_{u}}{\partial\operatorname{lift}\left(\mathbf{X}_{v}\right)}\operatorname{lift}\\ &=\left(\mathbf{I}_{c\times c},\mathbf{0}_{c\times c},\ldots,\mathbf{I}_{c\times c},\mathbf{0}_{c\times c}\right)\mathcal{H}(t,u,v)\mathbf{O}_{u}^{T}\mathbf{W}\mathbf{O}_{v}\left(\mathbf{I}_{c\times c},\mathbf{0}_{c\times c},\ldots,\mathbf{I}_{c\times c},\mathbf{0}_{c\times c}\right)^{T}\\ &=\mathcal{H}(t,u,v)\sum_{1\leq k_{1},\ k_{2}\leq k}\left(\mathbf{I}_{c\times c},\mathbf{0}_{c\times c}\right)\mathbf{O}_{u}^{T}\mathbf{W}\mathbf{O}_{v}\left(\mathbf{I}_{c\times c},\mathbf{0}_{c\times c}\right)^{T}\end{split}$
@@
We proceed by partitioning the indexing by $(k_{1},\ k_{2})$ into four cases. The first case is $C_{1}=\left\{(k_{1},k_{2})\text{ such that }\left(k_{1}\neq u,v\text{ and }k_{2}\neq u,v\right)\text{or}(k_{1}=k_{2}=u\text{ or }k_{1}=k_{2}=v)\right\}$ for which both $\mathbf{O}_{u}$ and $\mathbf{O}_{v}$ act like the identity. The second case is $C_{2}=\left\{(k_{1},k_{2})\text{ such that }k_{1}=u\text{ and }k_{2}\neq u,v\right\}$ where $\mathbf{O}_{u}$ flips the first $c$ rows with the second $c$ rows and $\mathbf{O}_{v}$ acts like the identity. $C_{3}=\left\{(k_{1},k_{2})\text{ such that }k_{2}=j\text{ and }k_{1}\neq u,v\right\}$ where $\mathbf{O}_{v}$ flips the first $c$ columns with the second $c$ columns, and $\mathbf{O}_{u}$ acts like the identity on the rows. Finally, the last case is when $k_{1}=u$ and $k_{v}=v$ in which $\mathbf{O}_{v}$ flips the rows and $\mathbf{O}_{v}$ flips the columns.
@@
$\begin{split}\ldots&=\mathcal{H}(t,u,v)\sum_{1\leq k_{1},\ k_{2}\leq k}\left(\mathbf{I}_{c\times c},\mathbf{0}_{c\times c}\right)\mathbf{O}_{u}^{T}\mathbf{W}\mathbf{O}_{v}\left(\mathbf{I}_{c\times c},\mathbf{0}_{c\times c}\right)^{T}\\ &=\mathcal{H}(t,u,v)\left(\mathbf{I}_{c\times c},\mathbf{0}_{c\times c}\right)\left[\sum_{(k_{1},k_{2})\in C_{1}}\begin{pmatrix}\mathbf{W}_{00}^{k_{1}k_{2}}&\mathbf{W}_{01}^{k_{1}k_{2}}\\ \mathbf{W}_{10}^{k_{1}k_{2}}&\mathbf{W}_{11}^{k_{1}k_{2}}\\ \end{pmatrix}+\sum_{(k_{1},k_{2})\in C_{2}}\begin{pmatrix}\mathbf{W}_{10}^{k_{1}k_{2}}&\mathbf{W}_{11}^{k_{1}k_{2}}\\ \mathbf{W}_{00}^{k_{1}k_{2}}&\mathbf{W}_{01}^{k_{1}k_{2}}\\ \end{pmatrix}\right.\\ &\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \left.+\sum_{(k_{1},k_{2})\in C_{3}}\begin{pmatrix}\mathbf{W}_{01}^{k_{1}k_{2}}&\mathbf{W}_{00}^{k_{1}k_{2}}\\ \mathbf{W}_{11}^{k_{1}k_{2}}&\mathbf{W}_{10}^{k_{1}k_{2}}\\ \end{pmatrix}+\begin{pmatrix}\mathbf{W}_{11}^{ij}&\mathbf{W}_{10}^{ij}\\ \mathbf{W}_{01}^{ij}&\mathbf{W}_{00}^{ij}\\ \end{pmatrix}\right]\left(\mathbf{I}_{c\times c},\mathbf{0}_{c\times c}\right)^{T}\end{split}$
@@
Finally, after applying $\left(\mathbf{I}_{c\times c},\mathbf{0}_{c\times c}\right)$ on the left and $\left(\mathbf{I}_{c\times c},\mathbf{0}_{c\times c}\right)^{T}$ on the right (an operation that selects the upper left $c\times c$ block), we observe that setting all $\mathbf{W}_{00}^{k_{1}k_{2}}=\mathbf{W}_{01}^{k_{1}k_{2}}=\mathbf{W}_{10}^{k_{1}k_{2}}$ to $\mathbf{0}_{c\times c}$ and setting $\mathbf{W}_{11}^{uv}:=\frac{1}{\mathcal{H}(t,u,v)}\frac{\partial{\left(L\mathbf{X}\right)_{u}}}{\partial\mathbf{x}_{v}}$ if the nodes corresponding to $k_{1}$ and $k_{2}$ lie in the same graph and $\mathbf{0}_{c\times c}$ otherwise. This allows us to conclude that any linear layer can be parameterized. ∎
Appendix D Discussion on compact uniform approximation versus uniform approximation
@@
A strong definition of expressivity that deals with infinite collections of graphs was proposed in Rosenbluth et al. [27]. This definition subsumes graph-isomorphism testing (where the input feature on graphs is constant) and is finer than graph-level function approximation since it works at the node level. Furthermore, it also deals with infinite families of graphs, as opposed to most mainstream theorems of GNN expressivity, which are proved for graphs of bounded size (e.g. Azizian and marc lelarge [24], Geerts and Reutter [25]) (see Section 2 for the notation and definition of features transformations).
Definition D.1 (From Rosenbluth et al. [27]).
@@
Note that this definition differs from our Definition 5.1 in that it requires uniform approximation over all graphs in $\mathcal{G}$ simultaneously, while we allow the width to vary with the finite subset $\mathcal{K}\subseteq\mathcal{G}$, similar to how classical results allow the width to vary with the compact set over which to approximate the function. Such a definition has proven useful in Rosenbluth et al. [27] to distinguish different aggregation functions and in Rosenbluth et al. [28] to distinguish MPNNs with virtual nodes from Transformers. However, we argue that the definition above is too strong for a finite parameter GNN. This is because it requires uniform approximation over a non-compact set, which contrasts with traditional work on expressivity and is generally unfeasible and impractical. Indeed, finite-parameters MLPs are not universal over the whole domain $\mathbb{R}$ under the $\ell_{\infty}$-norm. On an infinite collection of featured graphs, the topology is the disjoint union topology on $\bigsqcup_{\mathsf{G}\in\mathcal{G}}\mathbb{R}^{n_{\mathsf{G}}d}$, a compact subset consists of a finite set of graphs, and for each graph $\mathsf{G}$ only non-zero on a compact subset of $\mathbb{R}^{nd}$. For these reasons, we introduce Definition 5.1, which is still rich enough to distinguish between BuNNs and MPNNs.
@@ Appendix E Algorithmic and Implementation details
Computing the bundle maps. In our experiments, we noticed that having several bundles of dimension $2$ was more efficient than one bundle of large dimensions, while there was no clear performance gain when using higher dimensional bundles. To compute the $b$ bundle maps $\mathbf{O}_{v}$ we therefore only need $b$ rotation angles $\theta_{v}$, one per bundle. In our experiments, we use Housholder reflections using the python package noa [40] or direct parameterization. For direct parameterization, we do the following: since the matrix group $O(2)$ is disconnected, we always take $b$ to be even and parameterize half the bundles as rotation matrices $r(\theta)=\begin{pmatrix}\cos\left(\theta\right)&\sin\left(\theta\right)\\ -\sin\left(\theta\right)&\cos\left(\theta\right)\end{pmatrix}$ and the other half to correspond to matrices with determinant $-1$, which can be parameterized by $r^{*}(\theta)=\begin{pmatrix}\cos\left(\theta\right)&\sin\left(\theta\right)\\ \sin\left(\theta\right)&-\cos\left(\theta\right)\end{pmatrix}$. We compute the angles $\theta$ as in Equation 1 where the network $\boldsymbol{\phi}^{\left(\ell\right)}$ is either an MLP or a GNN. The network $\boldsymbol{\phi}$ is either shared across layers or differing at every layer.
@@
E.2 BuNN-Hop extension.
In practice we also explored a variation of BuNNs inspired by MixHop [34] which we call BuNN-Hop (described in Algorithm 3). We found that a slightly different model performed better on the Peptides benchmark datasets. One layer of the model can be derived by running a discrete time diffusion process with update $\mathbf{X}(t)=\mathbf{A}_{\mathcal{B}}\mathbf{X}(t-1)$ and integrating over it using attention term $\boldsymbol{\alpha}(t)$. The resulting update equation is $\mathbf{Y}=\sum_{k=1}^{K}\boldsymbol{\alpha}_{k}\mathbf{A}^{k}_{\mathcal{B}}\mathbf{X}$ where $\boldsymbol{\alpha}$ satisfy $\sum_{k=1}^{K}\boldsymbol{\alpha}_{k}=1$.
1:Input: Normalized graph adjacency $\mathbf{A}$, Orthogonal maps $\mathbf{O}_{v}^{\left(\ell\right)}\ \forall v\in\mathsf{G}$, Node features $\mathbf{X}^{\left(\ell\right)}\in\mathbb{R}^{n\times d}$, Maximum degree $K$, Channel mixing matrix $\mathbf{W}^{\left(\ell\right)}$, bias $\mathbf{b}^{\left(\ell\right)}$, attention terms $\boldsymbol{\alpha}^{(\ell)}$
2:Output: Updated node features $\mathbf{X}^{\left(\ell\right)}$
3:$\mathbf{h}_{v}^{\left(\ell\right)}\leftarrow{\mathbf{O}_{v}^{\left(\ell\right)}}^{T}\mathbf{x}_{v}^{\left(\ell\right)}\ \ \forall v\in\mathsf{V}$triangleright Go to global representation
6:for $k=1,\ldots K$ do
7:     $\mathbf{H}^{\left(\ell\right)}\leftarrow\mathbf{A}\mathbf{H}^{\left(\ell\right)}$triangleright Term of degree $k$
8:     $\mathbf{X}^{\left(\ell+1\right)}\leftarrow\mathbf{X}^{\left(\ell+1\right)}+\boldsymbol{\alpha}^{(\ell)}_{k}\mathbf{H}^{\left(\ell\right)}$triangleright Approximation of degree $k$
9:end for
10:$\mathbf{x}_{v}^{\left(\ell+1\right)}\leftarrow{\mathbf{O}_{v}^{\left(\ell\right)}}^{T}\mathbf{x}_{v}^{\left(\ell+1\right)}\ \ \forall v\in\mathsf{V}$triangleright Return to local representation
11:return $\mathbf{X}^{\left(\ell+1\right)}$
Algorithm 3 BuNN-Hop layer implementation.

### v2

@@ Appendix B Proofs
Defining the ‘universal bundle’: We fix $k=\sum_{\mathsf{G}\in\mathcal{K}}\left|\mathsf{V}_{\mathsf{G}}\right|$, so we can interpret our embedding space as a lookup table where each index corresponds to a node $v\in\bigsqcup_{\mathsf{G}\in\mathcal{K}}V_{\mathsf{G}}$. In turn, we can think of the parameter matrix $\mathbf{W}\in\mathbb{R}^{\left(\sum_{G\in\mathcal{K}}\left|\mathsf{V}_{\mathsf{G}}\right|\right)\times\left(\sum_{G\in\mathcal{K}}\left|\mathsf{V}_{\mathsf{G}}\right|\right)}$ as a lookup table where each entry corresponds to a pair of nodes in our dataset $\mathcal{K}$. Still thinking of the indices of the $2ck$ dimensions as $2c$-dimensional vectors indexed by the $k$ nodes in our dataset, we define $\mathbf{O}_{u}\in O\left(2ck\right)$ as a block diagonal matrix with $k$ different $2c$-dimensional blocks, where the $k_{i}$th block is denoted $\mathbf{O}_{u}^{k_{i}}$. These are all set to the identity except for the block at the index corresponding to node $u$, which is defined as $\begin{pmatrix}\mathbf{\mathbf{0}_{c\times c}}&\mathbf{I}_{c\times c}\\ \mathbf{I}_{c\times c}&\mathbf{\mathbf{0}_{c\times c}}\end{pmatrix}$ which is a $2c\times 2c$ matrix that acts by permuting the first $c$ dimensions with the second $c$ dimensions.
@@
$\begin{split}&\frac{\partial\left(\operatorname{BuNN}\left(\mathbf{X}\right)\right)_{u}}{\partial\mathbf{x}_{v}}\\ &=\operatorname{pool}\frac{\partial\left(\widehat{\operatorname{BuNN}}\left(\operatorname{lift}\left(\mathbf{X}\right)\right)\right)_{u}}{\partial\operatorname{lift}\left(\mathbf{X}_{v}\right)}\operatorname{lift}\\ &=\left(\mathbf{I}_{c\times c},\mathbf{0}_{c\times c},\ldots,\mathbf{I}_{c\times c},\mathbf{0}_{c\times c}\right)\mathcal{H}(t,u,v)\mathbf{O}_{u}^{T}\mathbf{W}\mathbf{O}_{v}\left(\mathbf{I}_{c\times c},\mathbf{0}_{c\times c},\ldots,\mathbf{I}_{c\times c},\mathbf{0}_{c\times c}\right)^{T}\\ &=\mathcal{H}(t,u,v)\sum_{1\leq k_{1},\ k_{2}\leq k}\left(\mathbf{I}_{c\times c},\mathbf{0}_{c\times c}\right){\mathbf{O}_{u}^{k_{1}}}^{T}\mathbf{W}^{k_{1}k_{2}}\mathbf{O}^{k_{2}}_{v}\left(\mathbf{I}_{c\times c},\mathbf{0}_{c\times c}\right)^{T}\end{split}$
@@
We proceed by partitioning the indexing by $(k_{1},\ k_{2})$ into four cases. The first case is $C_{1}=\left\{(k_{1},k_{2})\text{ such that }\left(k_{1}\neq u,v\text{ and }k_{2}\neq u,v\right)\right\}$ for which both $\mathbf{O}^{k_{1}}_{u}$ and $\mathbf{O}^{k_{2}}_{v}$ act like the identity. The second case is $C_{2}=\left\{(k_{1},k_{2})\text{ such that }k_{1}=u\text{ and }k_{2}\neq v\right\}$ where $\mathbf{O}^{k_{1}}_{u}$ flips the first $c$ rows with the second $c$ rows and $\mathbf{O}^{k_{2}}_{v}$ acts like the identity. $C_{3}=\left\{(k_{1},k_{2})\text{ such that }k_{2}=v\text{ and }k_{1}\neq u\right\}$ where $\mathbf{O}^{k_{2}}_{v}$ flips the first $c$ columns with the second $c$ columns, and $\mathbf{O}^{k_{1}}_{u}$ acts like the identity on the rows. Finally, the last case is when $k_{1}=u$ and $k_{2}=v$ in which $\mathbf{O}^{k_{1}}_{u}$ flips the rows, and $\mathbf{O}^{k_{1}}_{v}$ flips the columns. Note that by the injectivity assumption as defined in Definition D.1, each node has a unique positional encoding and therefore there such matrices can be parameterized for all nodes simultaneously.
@@
$\begin{split}\ldots&=\mathcal{H}(t,u,v)\sum_{1\leq k_{1},\ k_{2}\leq k}\left(\mathbf{I}_{c\times c},\mathbf{0}_{c\times c}\right){\mathbf{O}^{k_{1}}_{u}}^{T}\mathbf{W}^{k_{1}k_{2}}\mathbf{O}^{k_{2}}_{v}\left(\mathbf{I}_{c\times c},\mathbf{0}_{c\times c}\right)^{T}\\ &=\mathcal{H}(t,u,v)\left(\mathbf{I}_{c\times c},\mathbf{0}_{c\times c}\right)\left[\sum_{(k_{1},k_{2})\in C_{1}}\begin{pmatrix}\mathbf{W}_{00}^{k_{1}k_{2}}&\mathbf{W}_{01}^{k_{1}k_{2}}\\ \mathbf{W}_{10}^{k_{1}k_{2}}&\mathbf{W}_{11}^{k_{1}k_{2}}\\ \end{pmatrix}+\sum_{(k_{1},k_{2})\in C_{2}}\begin{pmatrix}\mathbf{W}_{10}^{k_{1}k_{2}}&\mathbf{W}_{11}^{k_{1}k_{2}}\\ \mathbf{W}_{00}^{k_{1}k_{2}}&\mathbf{W}_{01}^{k_{1}k_{2}}\\ \end{pmatrix}\right.\\ &\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \left.+\sum_{(k_{1},k_{2})\in C_{3}}\begin{pmatrix}\mathbf{W}_{01}^{k_{1}k_{2}}&\mathbf{W}_{00}^{k_{1}k_{2}}\\ \mathbf{W}_{11}^{k_{1}k_{2}}&\mathbf{W}_{10}^{k_{1}k_{2}}\\ \end{pmatrix}+\begin{pmatrix}\mathbf{W}_{11}^{uv}&\mathbf{W}_{10}^{uv}\\ \mathbf{W}_{01}^{uv}&\mathbf{W}_{00}^{uv}\\ \end{pmatrix}\right]\left(\mathbf{I}_{c\times c},\mathbf{0}_{c\times c}\right)^{T}\\ &=\mathcal{H}(t,u,v)\left[\sum_{(k_{1},k_{2})\in C_{1}}\mathbf{W}_{00}^{k_{1}k_{2}}+\sum_{(k_{1},k_{2})\in C_{2}}\mathbf{W}_{10}^{k_{1}k_{2}}+\sum_{(k_{1},k_{2})\in C_{3}}\mathbf{W}_{01}^{k_{1}k_{2}}+\mathbf{W}_{11}^{uv}\right]\end{split}$
@@
Where the last line is obtained by applying $\left(\mathbf{I}_{c\times c},\mathbf{0}_{c\times c}\right)$ on the left and $\left(\mathbf{I}_{c\times c},\mathbf{0}_{c\times c}\right)^{T}$ on the right, an operation that selects the upper left $c\times c$ block. We observe that setting all $\mathbf{W}_{00}^{k_{1}k_{2}}=\mathbf{W}_{01}^{k_{1}k_{2}}=\mathbf{W}_{10}^{k_{1}k_{2}}$ to $\mathbf{0}_{c\times c}$ and setting $\mathbf{W}_{11}^{uv}:=\frac{1}{\mathcal{H}(t,u,v)}\frac{\partial{\left(L\mathbf{X}\right)_{u}}}{\partial\mathbf{x}_{v}}$ if the nodes corresponding to $u$ and $v$ lie in the same graph and $\mathbf{0}_{c\times c}$ otherwise. This allows us to conclude that any linear layer can be parameterized, completing the proof of the theorem. ∎
Appendix C Discussion on compact uniform approximation versus uniform approximation
@@
A strong definition of expressivity that deals with infinite collections of graphs was proposed in Rosenbluth et al. (2023). This definition subsumes graph-isomorphism testing (where the input feature on graphs is constant). Furthermore, it also deals with infinite families of graphs, as opposed to most mainstream theorems of GNN expressivity, which are proved for graphs of bounded size (e.g. Azizian & Lelarge (2021); Geerts & Reutter (2022)). See Section 2 for the notation and definition of features transformations.
Definition C.1 (From Rosenbluth et al. (2023)).
@@
Note that this definition differs from our Definition 5.1 in that it requires uniform approximation over all graphs in $\mathcal{G}$ simultaneously, while we allow the width to vary with the finite subset $\mathcal{K}\subseteq\mathcal{G}$, similar to how classical results allow the width to vary with the compact set over which to approximate the function. Such a definition has proven useful in Rosenbluth et al. (2023) to distinguish different aggregation functions and in Rosenbluth et al. (2024) to distinguish MPNNs with virtual nodes from Transformers. However, we argue that the definition above is too strong for a finite parameter GNN. This is because it requires uniform approximation over a non-compact set, which contrasts with traditional work on expressivity and is generally unfeasible and impractical. Indeed, finite-parameters MLPs are not universal over the whole domain $\mathbb{R}$ under the $\ell_{\infty}$-norm. On an infinite collection of featured graphs, the topology is the disjoint union topology on $\bigsqcup_{\mathsf{G}\in\mathcal{G}}\mathbb{R}^{n_{\mathsf{G}}d}$, a compact subset consists of a finite set of graphs, and for each graph $\mathsf{G}$ only non-zero on a compact subset of $\mathbb{R}^{nd}$. For these reasons, we introduce Definition 5.1, which is still rich enough to distinguish between BuNNs and MPNNs.
Appendix D Why classical arguments do not apply to compact uniform approximation
A seminal result in GNN expressivity is Theorem 2 in Morris et al. (2019). In this section, we discuss why the arguments do not hold for compact uniform approximation, even when enriched with injective positional encodings. We start by formally defining injective positional encodings.
Definition D.1.
A positional encoding $\pi$ on a graph $\mathsf{G}$ is a map $\pi(\mathsf{G}):\mathsf{V}\rightarrow\mathbb{R}^{k}$ which assigns a $k$-dimensional feature to every node . A positional encoding $\pi$ on a family of graphs $\mathcal{G}$ is a positional encoding on all graphs $\mathsf{G}\in\mathcal{G}$. A positional encoding $\pi$ is injective on $\mathcal{G}$ for any graph $\mathsf{G}\in\mathcal{G}$ and every node $u\in\mathsf{G}$ there is no other node $v\in\mathsf{H}$ for any $\mathsf{H}\in\mathcal{G}$ with $\pi(u)=\pi(v)$. In other words, each positional encoding corresponds to a unique node in the dataset. [^4] Note that on symmetric graphs such as the cycle, it is not possible to have injectivity if the positional encoding is equivariant. However, in practice many steps of the $1$-WL color refinement will be equivariant and satisfy such assumption on most graphs.
The setting in Morris et al. (2019) Theorem $2$ considers a single graph with fixed node features (or colors/labels). In contrast, our Theorem is for a family of graphs, and more importantly, for each graph $\mathsf{G}$, the node features are not fixed but can vary in any compact subspace of feature space. This means that for a single graph of size $n$, our statement holds, for example, on the unit cube $[0,1]^{n}$, while the result in Morris et al. (2019) only holds for a specific point in $\mathbb{R}^{n}$.
Fixing the node features is precisely what makes the construction in Morris et al. (2019) possible. Indeed the proof starts by assuming that the initial coloring is “linearly independent modulo equality”, denoted by $\mathbf{F}^{(0)}_{l,0}$ in their proof. This property on node features is indeed central to the construction. It is used several times, for example “Observe that colors are represented by linearly independent row vectors” and ”$\mathbf{F}^{(t+1)}_{l,0}$ is linearly independent modulo equality”. Such an assumption is possible when dealing with fixed node features: since there are at most $n$ colors, it suffices to take a one-hot encoding of those colors, which fits in a feature space of dimension at most $n$. This assumption is also crucial for the injectivity of the sum aggregation.
[…]
@@ Appendix E Algorithmic and Implementation details
Computing the bundle maps. In our experiments, we noticed that having several bundles of dimension $2$ was more efficient than one bundle of large dimensions, while there was no clear performance gain when using higher dimensional bundles. To compute the $b$ bundle maps $\mathbf{O}_{v}$ we therefore only need $b$ rotation angles $\theta_{v}$, one per bundle. In our experiments, we use Housholder reflections using the python package Obukhov (2021) or direct parameterization. For direct parameterization, we do the following: since the matrix group $O(2)$ is disconnected, we always take $b$ to be even and parameterize half the bundles as rotation matrices $r(\theta)=\begin{pmatrix}\cos\left(\theta\right)&\sin\left(\theta\right)\\ -\sin\left(\theta\right)&\cos\left(\theta\right)\end{pmatrix}$ and the other half to correspond to matrices with determinant $-1$, which can be parameterized by $r^{*}(\theta)=\begin{pmatrix}\cos\left(\theta\right)&\sin\left(\theta\right)\\ \sin\left(\theta\right)&-\cos\left(\theta\right)\end{pmatrix}$. We compute the angles $\theta$ as in Equation 1 where the network $\bm{\phi}^{\left(\ell\right)}$ is either an MLP or a GNN. The network $\bm{\phi}$ is either shared across layers or differing at every layer.